OpenStudy (anonymous):

write sin(34 degrees) as a cofunction of a complementary angle

OpenStudy (anonymous):

hmm... cofunction... sin(angle)=cos(90º-angle) ? so sin(34º)=cos(90º-34º)=cos(56º) not sure... XP

OpenStudy (anonymous):

right. that's what i was thinking, but it seems so simple!